# SVEMnet: An R package for Self-Validated Elastic-Net Ensembles and Multi-Response Optimization in Small-Sample Mixture–Process Experiments

### v2

@@ 3 Example LNP Application
We illustrate the workflow using the bundled simulated LNP formulation screening data, which include four mixture components (PEG, Helper, Ionizable, Cholesterol), a categorical ionizable lipid type, two continuous process factors (N_P_ratio and flow_rate), a categorical Operator factor, and three responses: Potency, Size, and PDI. This example parallels the SVEM-based mixture–process LNP workflow of Karl et al. [17].
@@ 5 Ionizable_Lipid_Type + N_P_ratio + flow_rate,
11)
12form_pot <- bigexp_formula(spec, "Potency")
13form_siz <- bigexp_formula(spec, "Size")
14form_pdi <- bigexp_formula(spec, "PDI")
@@ 5 PDI = form_pdi),
Table 1: Permutation whole-model test results for the three responses (SVEMnet). Values are permutation-based and subject to Monte Carlo error.
Response Approx. whole-model $p$-value
Size $1.00\text{\times}{10}^{-16}$
Potency $1.35\text{\times}{10}^{-14}$
PDI 0.72
The approximate whole-model $p$-values are summarized in Table 1: Size and Potency show strong global signal, whereas PDI does not (although lack of signal in this heuristic alone should not be used to conclude that there is no relationship between the study factors and PDI [5]). Figure 2 shows the associated whole-model significance diagnostic, generated by svem_wmt_multi() under the shared expansion and mixture constraints. Clear separation between the “Original” and permutation-based distance distributions indicates strong global factor signal, as in Figure 3 of Karl [5].
Figure 2: Whole-model significance diagnostic for Potency, Size, and PDI: distributions of Mahalanobis-like distances for the original data versus permuted replicates under the shared deterministic expansion and mixture constraints; cf. Figure 3 in [5].
@@ 3 Size = list(upper = 100),
Analogous calls with target = "uncertainty_measure" and target = "p_joint_mean" (and score_table set to either the random-search candidates or the original screening runs) were used to obtain the candidates summarized in Tables 2 and 3. The replication script contains the full set of svem_select_from_score_table() calls [40].
In routine use these tools are embedded in a simple sequential loop: fit SVEM on the current data under a fixed deterministic expansion; use svem_score_random() and svem_select_from_score_table() to generate optimal, exploration, and mean-in-spec medoids; run those settings and append the new rows; then refit and repeat. Candidate tables for laboratory use can be exported with svem_export_candidates_csv(); the Supplementary CSV file was generated using this function and is included in the Mendeley Data archive [40].
Tables 2 and 3 show four SVEM candidates from the simulated first-round lipid screening. The first row corresponds to the best existing screened run by multi-response score and provides a baseline. This row displays the original level of the blocking factor, Operator, which may differ from the level shown for the random-search candidates; the latter are evaluated at the reference Operator level used for scoring. The additive offsets for the Operator levels are visible in the fitted coefficients (e.g., coef(fit_pot)), and predictions for other Operator levels can be obtained via predict(fit_pot, newdata = ...). The score-optimal run from the random search trades a modest reduction in predicted Potency for a large improvement in Size, yielding an overall desirability of one. In this example, the wmt_scorewmt\_score optimum coincides with the score-optimal run, so no additional row is shown. The “In-spec” optimum illustrates a candidate that lies comfortably inside the specification region. The exploration target has the highest uncertainty measure $U(x)$ and sits near the edge of the feasible mixture region, with poor predicted Potency.
Table 2: Representative SVEM candidate recipes from the lipid screening example. Mixture components are proportions on the 0–1 scale; process factors are in their native units. All random-search candidates arbitrarily set Operator to A.
@@
Best screened 0.047 0.430 0.140 0.383 H101 6.0 2.5 B
Max Score 0.048 0.596 0.101 0.256 H102 11.9 2.9 A
In-spec 0.040 0.415 0.193 0.352 H101 9.2 2.5 A
Exploration 0.012 0.593 0.209 0.185 H103 6.3 2.7 A
Table 3: Predicted behavior for the candidate recipes in Table 2. Responses are SVEM point predictions. The columns score and wmt_score are scalar desirability measures, $p_{\mathrm{joint,mean}}$ is the estimated probability that the mean responses satisfy the specification limits (Potency $\geq 78$, Size $\leq 100$, PDI $\leq 0.25$), and uncert. is the dimensionless uncertainty measure $U(x)$.
Scenario Potency Size PDI score wmt_score $p_{\mathrm{joint,mean}}$ uncert.
Best screened 91.5 88.2 0.21 0.83 0.91 0.93 0.77
Max Score 89.4 47.7 0.17 1.00 1.00 1.00 0.50
In-spec 86.3 89.8 0.20 0.81 0.83 1.00 0.18
Exploration 71.0 49.4 0.21 0.29 0.29 0.05 0.92
@@ 3.3 Computational environment and runtime
All timings were obtained on a Windows 11 x64 desktop with an Intel Core Ultra 7 265 CPU (20 cores, 2.4 GHz) and 64 GB RAM, running R 4.5.2 with SVEMnet 3.1.4 and glmnet 4.1-10 and parallel execution enabled for the WMT. Users should note that on Windows, the parallelization of the WMT will trigger a one-time Windows Firewall prompt for R when worker processes are created; this is due to local socket connections between R processes and does not involve external network access. The user may click Cancel on this alert window and the program will still run in parallel.
@@ 4 Simulation study
We carried out simulations comparing SVEMnet to repeated cross-validated elastic-net baselines over grids of sample sizes and signal strengths, where “signal strength” refers to the proportion of variance in the data explained by the underlying model (target $R^{2}$). The baselines use a thin wrapper glmnet_with_cv() around cv.glmnet, matching the $\alpha$ grid and penalty path to the SVEMnet settings. For each configuration we ran both the default two-point grid $\alpha\in\{0.5,1\}$ and a lasso-only grid $\alpha=1$, which behaved indistinguishably in these simulations. All simulation scripts and CSV result files are provided in the replication bundle of Karl [40].
@@ 4.1 Design and metrics
Each replicate uses a sparse quadratic response surface in four continuous factors $X_{1}{:}X_{4}\in[-1,1]$ and one three-level factor $X_{5}\in{\mathrm{L1},\mathrm{L2},\mathrm{L3}}$ (sum-to-zero coding). Training data are Latin hypercube samples for $X_{1}{:}X_{4}$ with balanced $X_{5}$ levels [32], with $n_{\text{total}}\in\{15,20,\dots,50\}$. As noted in the introduction, we use Latin hypercubes here for simulation convenience; in practice, more sophisticated optimal or space-filling designs are typically preferable when constructing the initial experimental design [41, 28, 30, 31, 29].
The quadratic model has $p_{\mathrm{full}}=25$ coefficients (including the intercept), so $n_{\text{total}}=p_{\mathrm{full}}=25$ is the interpolation boundary. To induce sparsity, in each simulation replicate and for each coefficient $j$, draw $\pi_{j}\sim\mathrm{Beta}(0.5,0.5)$, $Z_{j}\sim\mathrm{Bernoulli}(\pi_{j})$, and $E_{j}$ from a mean-zero Laplace distribution, and set $\beta_{j}=Z_{j}E_{j}$. This produces many exact zeros with Laplace-distributed nonzero effects.
For the Gaussian case, we evaluate the noiseless surface on a large grid to estimate its standard deviation $\sigma_{f}$, add Gaussian noise to the training responses so that the expected domain $R^{2}$ matches a target $R^{2}\in\{0.5,0.9\}$, and score all methods on a separate noiseless 10 000-point holdout. Performance is summarized by log-normalized RMSE (log-NRMSE), where RMSE on the holdout is normalized by the empirical standard deviation of the noiseless surface on that grid.
@@
To reduce Monte Carlo error, the summary curves in Figures 3, 4, and 5 average over 1000 independent simulation replicates for each combination of $n_{\text{total}}$, target $R^{2}$, model order, and fitting method (Setting). In these overview plots we fix a common set of tuning defaults. For Gaussian responses, cv.glmnet uses standard (non-relaxed) elastic-net paths (relax = FALSE) and SVEMnet uses relaxed base learners (relax = TRUE); post-hoc debiasing is disabled for all methods and the elastic-net mixing grid is held at $\alpha\in\{0.5,1\}$. For binomial responses, all methods use non-relaxed paths (relax = FALSE) with the same default $\alpha$ grid. These settings are used for all model orders shown in the figures.
@@ 4.2 Gaussian results
Figure 3: Effect of expansion order on mean log-NRMSE (lower is better) plotted against $n_{\text{total}}$ for the Gaussian simulations. Rows correspond to target $R^{2}\in\{0.5,0.9\}$ and columns to the model order used in fitting. Colors and line types indicate the selector: repeated-cv.glmnet (blue dashed), wAIC (orange dotted), wBIC (green solid), and the loss-only wSSE selector (purple dot–dash). Light gray vertical lines in the middle and right-hand panels mark the interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$ for Orders 2 and 3 ($p_{\mathrm{full}}=25$ and $p_{\mathrm{full}}=45$, respectively). Simulation settings match Figure 3: non-relaxed cv.glmnet, relaxed SVEMnet, no debiasing. The wSSE selector exhibits a pronounced spike in holdout log-NRMSE near $n_{\text{total}}=p_{\mathrm{full}}$, whereas wAIC and wBIC remain stable across the interpolation boundary.
Figure 4: Mean median-selected size $k_{\lambda}$ versus $n_{\text{total}}$ for the Gaussian simulations, stratified by target $R^{2}$ (rows) and model order (columns). Colors and line types indicate the selector: repeated-cv.glmnet (blue dashed), wAIC (orange dotted), wBIC (green solid), and the loss-only wSSE selector (purple dot–dash). The light gray vertical lines in the middle and right-hand panels mark the interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$ for Orders 2 and 3 ($p_{\mathrm{full}}=25$ and $p_{\mathrm{full}}=45$, respectively). Simulation settings match Figure 3: non-relaxed cv.glmnet, relaxed SVEMnet, no debiasing. The wSSE selector exhibits a pronounced spike in $k_{\lambda}$ near $n_{\text{total}}=p_{\mathrm{full}}$.
@@
    Information criteria versus loss-only selectors. Across the grid, SVEMnet with validation-weighted information criteria (wAIC, wBIC) dominates the historical loss-only wSSE selector in terms of mean log-NRMSE on the noiseless holdout. wAIC often slightly improves on the repeated cv.glmnet for Order 2 and 3. The best overall configuration in these simulations is SVEM-wAIC with relaxed elastic-net base learners and no debias step; wBIC is close behind (Figure 3 and Appendix Table LABEL:tab:tukey-hsd).
@@
    Relaxation and debiasing. For the SVEM information-criterion selectors, relaxed base learners improve performance relative to non-relaxed fits, while relaxation degrades cv.glmnet performance. The best-performing overall configuration in these simulations is SVEM-wAIC with relaxed base learners and no debias step; wBIC is close behind (Figure 3 and Table LABEL:tab:tukey-hsd).
@@
    Peaking near the interpolation boundary. For each simulation run we recorded $k_{\mathrm{median}}$, the median number of active parameters $k_{\lambda}$ across SVEM bootstrap replicates; Figure 4 shows the mean of these medians over simulation replicates. The wSSE selector shows a pronounced spike in both mean log-NRMSE (Figure 3) and median selected model size $k_{\lambda}$ (Figure 4) when $n_{\text{total}}$ equals the expansion dimensions: $p_{\mathrm{full}}=25$ for Order 2 and $p_{\mathrm{full}}=45$ for Order 3 ($p_{\mathrm{full}}=7$ for Order 1 is not included in the range of the graph). This is the “peaking” phenomenon near the interpolation boundary in small-sample regression [16]. In contrast, wAIC and wBIC remain smooth in both error and $k_{\lambda}$.
@@ 4.3 Binomial results
Figure 5: Binomial simulation summary (response LogLoss_Holdout versus true probabilities). Mean log-loss on a large holdout versus $n_{\text{total}}$, stratified by target $R^{2}$ (rows) and model order (columns). Colors and line types indicate the selector: repeated-cv.glmnet, wAIC, wBIC, and the loss-only wSSE selector. All methods use non-relaxed paths (relax = FALSE) with the default mixing grid $\alpha\in\{0.5,1\}$. Light gray vertical lines mark the interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$ for Orders 2 and 3 ($p_{\mathrm{full}}=25$ and $p_{\mathrm{full}}=45$), where wSSE shows a pronounced spike in log-loss.

### v3

@@ 3 Example LNP Application
We illustrate the workflow using the bundled simulated LNP formulation screening data, which include four mixture components (PEG, Helper, Ionizable, Cholesterol), a categorical ionizable lipid type, two discrete numeric process factors (N_P_ratio and flow_rate), a categorical Operator factor, and three responses: Potency, Size, and PDI. This example parallels the SVEM-based mixture–process LNP workflow of Karl et al. [19].
@@ 5 Ionizable_Lipid_Type + N_P_ratio + flow_rate,
12)
13form_pot <- bigexp_formula(spec, "Potency")
14form_siz <- bigexp_formula(spec, "Size")
15form_pdi <- bigexp_formula(spec, "PDI")
@@ 5 PDI = form_pdi),
The approximate whole-model $p$-values under the default permutation budget were Size $p\approx 1.0\times 10^{-16}$, Potency $p\approx 4.1\times 10^{-9}$, and PDI $p\approx 0.77$. Thus Size and Potency show strong global signal, whereas PDI does not (although lack of signal in this heuristic alone should not be used to conclude that there is no relationship between the study factors and PDI [5]). Figure 3 shows the associated whole-model significance diagnostic, generated by svem_wmt_multi() under the shared expansion and mixture constraints. Clear separation between the “Original” and permutation-based distance distributions indicates strong global factor signal, as in Figure 3 of Karl [5]. B.4 reports a simple stability check over random seeds and permutation budgets for this example.
Figure 3: Whole-model significance diagnostic (WMT) for Potency, Size, and PDI: distance distributions for the original data versus permuted fits under the shared expansion and mixture constraints.
@@ 3 Size = list(upper = 100),
Analogous calls with target = "uncertainty_measure" and target = "p_joint_mean" (and score_table set to either the random-search candidates or the original screening runs) were used to obtain the candidates summarized in Tables 1 and 2. The replication script contains the full set of svem_select_from_score_table() calls [44]. Candidate tables for laboratory use can be exported with svem_export_candidates_csv(); the Supplementary CSV file was generated using this function and is included in the Mendeley Data archive [44].
Tables 1 and 2 show four SVEM candidates from the simulated first-round lipid screening. The first row corresponds to the best existing screened run by multi-response score and provides a baseline. This row displays the original level of the blocking factor, Operator, which may differ from the level shown for the random-search candidates; the latter are evaluated at the reference Operator level used for scoring. The additive offsets for the Operator levels are visible in the fitted coefficients (e.g., coef(fit_pot)), and predictions for other Operator levels can be obtained via predict(fit_pot, newdata = ...). The score-optimal run from the random search trades a modest reduction in predicted Potency for a large improvement in Size, yielding an overall desirability of one. In this example, the wmt_score optimum coincides with the score-optimal run, so no additional row is shown. The “In-spec” optimum illustrates a candidate that lies comfortably inside the specification region. The exploration target has the highest uncertainty measure $U(x)$ and sits near the edge of the feasible mixture region, with poor predicted Potency. Table 2 reports the SVEM point predictions together with the primary desirability score (score), the WMT-reweighted diagnostic score (wmt_score), the mean-in-spec proxy $p_{\mathrm{joint,mean}}$, and the uncertainty measure $U(x)$ (Section 2.6).
Table 1: Representative SVEM candidate recipes from the lipid screening example. Mixture components are proportions on the 0–1 scale; process factors are in their native units. All random-search candidates are evaluated at the reference (modal) Operator level A.
@@
Best screened 0.047 0.430 0.140 0.383 H101 6 2.5 B
Max Score 0.048 0.583 0.110 0.259 H101 10 2.5 A
In-spec 0.040 0.458 0.165 0.337 H103 14 2.5 A
Exploration 0.011 0.592 0.123 0.273 H103 6 2.5 A
Table 2: Predicted behavior for the candidate recipes in Table 1. Responses are SVEM point predictions.
Scenario Potency Size PDI score wmt_score $p_{\mathrm{joint,mean}}$ $U(x)$
Best screened 91.5 88.2 0.21 0.83 0.88 0.93 0.77
Max Score 89.1 50.4 0.17 1.00 1.00 1.00 0.39
In-spec 84.9 76.5 0.17 0.83 0.82 1.00 0.73
Exploration 74.3 51.2 0.20 0.45 0.54 0.09 0.94
@@ 3.3 Computational environment and runtime
All timings were obtained on a Windows 11 x64 desktop with an Intel Core Ultra 7 265 CPU (20 cores, 2.4 GHz) and 64 GB RAM, running R 4.5.2 with SVEMnet 3.2.0 and glmnet 4.1-10 and parallel execution enabled for the WMT. Users should note that on Windows, parallelization of the WMT may trigger a one-time Windows Firewall prompt for R when worker processes are created; this is due to local socket connections between R processes and does not involve external network access. In typical setups the computation will still run if the prompt is dismissed.
@@ 4 Simulation study
We carried out simulations comparing SVEMnet to repeated cross-validated elastic-net baselines over grids of sample sizes and signal strengths, where “signal strength” refers to the proportion of variance in the data explained by the underlying model (target $R^{2}$). The baselines use a wrapper glmnet_with_cv() around cv.glmnet, matching the $\alpha$ grid and penalty path to the SVEMnet settings. For each configuration we ran both the default two-point grid $\alpha\in\{0.5,1\}$ and a lasso-only grid $\alpha=1$, which behaved indistinguishably in these simulations. All simulation scripts and CSV result files are provided in the replication bundle of Karl [44].
@@ 4.1 Design and metrics
Each replicate uses a sparse quadratic response surface in four continuous factors $X_{1}{:}X_{4}\in[-1,1]$ and one three-level factor $X_{5}\in{\mathrm{L1},\mathrm{L2},\mathrm{L3}}$ (sum-to-zero coding). Training data are Latin hypercube samples for $X_{1}{:}X_{4}$ with balanced $X_{5}$ levels [34], with $n_{\text{total}}\in\{15,20,\dots,50\}$. As noted in the introduction, we use Latin hypercubes here for simulation convenience; in practice, more sophisticated optimal or space-filling designs are typically preferable when constructing the initial experimental design [45, 30, 32, 33, 31]. B.2 reports representative robustness checks using a face-centered central composite design (CCD) in place of the Latin-hypercube baseline and using denser/alternative coefficient regimes. All simulations treat runs as independent (single error stratum); hierarchical split-plot dependence is outside the scope of SVEM and is not considered.
The quadratic model has $p_{\mathrm{full}}=25$ parameters (including the intercept), so $n_{\text{total}}=p_{\mathrm{full}}=25$ is the interpolation boundary. To induce sparsity, in each simulation replicate and for each coefficient $j$, draw $\pi_{j}\sim\mathrm{Beta}(0.5,0.5)$, $Z_{j}\sim\mathrm{Bernoulli}(\pi_{j})$, and $E_{j}$ from a mean-zero Laplace distribution, and set $\beta_{j}=Z_{j}E_{j}$. This produces many exact zeros with Laplace-distributed nonzero effects.
For the Gaussian case, we evaluate the noiseless response surface $f(x)$ on a large space-filling grid to estimate its domain-scale standard deviation $\sigma_{f}=\mathrm{sd}\{f(x)\}$. We then add Gaussian noise to the training responses using
$Y=f(x)+\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}),\qquad\sigma_{\varepsilon}=\sigma_{f}\sqrt{\frac{1-R^{2}}{R^{2}}},$
so that $\mathrm{Var}\{f(x)\}/\bigl(\mathrm{Var}\{f(x)\}+\mathrm{Var}(\varepsilon)\bigr)\approx R^{2}$ matches a target $R^{2}\in\{0.5,0.9\}$. All methods are scored on a separate noiseless 10,000-point holdout. Performance is summarized by the normalized RMSE, $\mathrm{NRMSE}=\mathrm{RMSE}/\mathrm{sd}\{f(x)\}$ (with $\mathrm{sd}\{f(x)\}$ computed on the holdout grid), and we report $\log_{e}(\mathrm{NRMSE})$.
@@
To reduce Monte Carlo error, the summary curves in Figures 4, 5, and 6 average over 1000 independent simulation replicates for each combination of $n_{\text{total}}$, target $R^{2}$, model order, and fitting method (Setting). In these overview plots we fix a common set of tuning defaults. For Gaussian responses, cv.glmnet uses standard (non-relaxed) elastic-net paths (relax = FALSE) and SVEMnet uses relaxed base learners (relax = TRUE); post-hoc debiasing is disabled for all methods and the elastic-net mixing grid is held at $\alpha\in\{0.5,1\}$. For binomial responses, all methods use non-relaxed paths (relax = FALSE) with the same default $\alpha$ grid. These settings are used for all model orders shown in the figures.
@@ 4.2 Gaussian results
Figure 4 summarizes mean log-NRMSE on a large noiseless holdout as a function of $n_{\text{total}}$ for target $R^{2}\in\{0.5,0.9\}$ (rows) and three right-hand-side expansion orders (columns). Curves compare repeated $5\timesCV$ elastic-net baselines (cv.glmnet) with SVEMnet selectors (wAIC, wBIC, and the historical loss-only wSSE). For these overview plots, cv.glmnet uses standard (non-relaxed) elastic-net paths, SVEMnet uses relaxed base learners for Gaussian responses, and post-hoc debiasing is disabled for all methods.
Figure 4: Gaussian simulations: mean holdout log-NRMSE (lower is better) versus $n_{\text{total}}$ by target $R^{2}$ (rows) and fitted expansion order (columns). Vertical lines mark interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$.
To connect predictive error to selected model complexity, Figure 5 reports the mean of the within-replicate median active-set size $k_{\lambda}$ (number of nonzero coefficients) selected across SVEM bootstrap replicates. The loss-only wSSE selector shows a pronounced spike near $n_{\text{total}}=p_{\mathrm{full}}$, whereas wAIC and wBIC remain stable.
Figure 5: Gaussian simulations: mean selected model size (median $k_{\lambda}$ across SVEM replicates) versus $n_{\text{total}}$ by target $R^{2}$ (rows) and expansion order (columns). Vertical lines mark $n_{\text{total}}=p_{\mathrm{full}}$.
@@
    Information criteria versus loss-only selectors. Across the grid, SVEMnet with validation-weighted information criteria (wAIC, wBIC) consistently outperforms the historical loss-only wSSE selector in mean holdout log-NRMSE. Relative to repeated cv.glmnet, SVEMnet-wAIC is typically best or tied for best for the correctly specified and mildly overspecified expansions (Orders 2 and 3), with wBIC close behind (Figure 4 and Appendix Table LABEL:tab:tukey-hsd).
@@
    Relaxation and debiasing. For SVEMnet under wAIC/wBIC, enabling relaxed base learners improves holdout performance in the Gaussian setting, whereas relaxation tends to degrade the repeated cv.glmnet baselines. The optional post-hoc debiasing step ($y\sim\hat{y}$ fit on the training data) does not improve mean holdout error in these simulations and is mildly unfavorable on average, so we recommend debias = FALSE as the default for Gaussian responses (Table LABEL:tab:tukey-hsd).
@@
    Peaking near the interpolation boundary. For each simulation run we recorded $k_{\mathrm{median}}$, the median number of active parameters $k_{\lambda}$ across SVEM bootstrap replicates; Figure 5 shows the mean of these medians over simulation replicates. The wSSE selector shows a pronounced spike in both mean log-NRMSE (Figure 4) and median selected model size $k_{\lambda}$ (Figure 5) when $n_{\text{total}}$ equals the expansion dimensions: $p_{\mathrm{full}}=25$ for Order 2 and $p_{\mathrm{full}}=45$ for Order 3 ($p_{\mathrm{full}}=7$ for Order 1 is not included in the range of the graph). This is the “peaking” phenomenon near the interpolation boundary in small-sample regression [16]. In contrast, wAIC and wBIC– with their effective sample size guardrails – remain smooth in both error and $k_{\lambda}$.
@@ 4.3 Binomial results
Figure 6 summarizes the corresponding mean holdout log-loss curves for the overview settings (all methods non-relaxed; default $\alpha$ grid).
Figure 6: Binomial simulations: mean holdout log-loss versus $n_{\text{total}}$ by target $R^{2}$ (rows) and fitted expansion order (columns). Vertical lines mark interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$.
@@ 5 Conclusion
SVEMnet is designed for single-error-stratum settings in which runs can be treated as independent resampling units. Hierarchical split-plot and other multi-stratum designs with random effects are not supported by SVEM and would require substantial methodological changes beyond the scope of this paper. In addition, the validation-weighted information-criterion analogs, WMT “$p$-values,” and mean-in-specification probabilities are intended as heuristic tuning, ranking, and diagnostic summaries rather than as fully calibrated inferential quantities. We therefore recommend reporting key tuning budgets and performing simple sensitivity checks (e.g., varying seeds, bootstrap/permutation budgets, desirability parameters, or expansion order) when conclusions are marginal.